OpenStudy (anonymous):

The ratio of the numerator to the denominator of a fraction is 2 to 3. If both the numerator and the denominator are increased by 2, the fraction becomes 3/4. What is the original fraction? Which of the following systems of equations can be used to solve the problem?

OpenStudy (anonymous):

Solve the following equations for x an y:\[\left\{\frac{x}{y}=\frac{2}{3},\frac{x+2}{y+2}=\frac{3}{4}\right\} \]

OpenStudy (anonymous):

I don't know how to do fractions

OpenStudy (anonymous):

If from the first equation, \[x=\frac{2 y}{3} \]Replace x in the second equation with 2y/3 and then solve for y. Can you proceed from this point?

OpenStudy (anonymous):

yes, thanks :)